The product of two postive numbers is 84. Find the two numbers if one number is 2 more than twice the other.
Answer by mananth(16946)

let one number be x
other number = 2x+2
..
x(2x+2)=84
2x^2+2x=84
/2
x^2+x=42
x^2+x-42=0
x+7x-6x-42=0
x(x+7)-6(x+7)=0
(x+7)(x-6)
x=6 one number
other number = 14